# Enhancing elusive clues in knowledge learning by contrasting attention of language models

### v1

@@ Introduction
Unlike previous work which was focus on distilling knowledge from large language models to make small language models perform better, our approach has improved the performance of large language models on not only synthetic , but also real-world datasets.
@@
We release the code and data used in this paper for reproducibility and further research111https://github.com/hushes-minutes/contrasting˙attention.
@@
The questions are formatted like “When was Liam Thompson born?”. Details on the training corpus and evaluation data are provided in the Appendix A.
@@
We show that even pretrained models with billions of parameters struggle to memorize facts perfectly in continual pretraining. Table 1 shows that while Gemma 2 (Team et al. 2024) and LLaMA 3 (Dubey et al. 2024) memorize the first two pieces of information (birth date and birth city) with high accuracy, they struggle to memorize the following three pieces of information (university, major, and company). This rules out the possibility that the performance deficiency is due to limited model size or insufficient pretraining.
@@
| Birth date | Birth city | University | Major | Company | ||
@@
We have shown that language models could achieve near-perfect accuracy in memorizing relationships that span a short distance in text, but struggle when they span a longer distance. In this section, we use attention weights as an interpretability tool to analyze the model’s behavior while learning long-range dependencies. We show that LLMs pay inadequately little attention to key information that is located further away, and more performant larger models can pay more attention to these information than smaller models.
@@
To visualize model’s attention weights when predicting the tail entities in a relationship, we extract the attention weights at the preposition tokens, i.e., the word immediately preceding the tail entity. For example, in the sentence “He received mentorship and guidance from faculty members at Sorbonne University”, the attention from the token “at” is extracted. Because the model is predicting the tail entity “Sorbonne University” at this position, the attention weights222To simplify analysis, we took the approach of averaging the attention weights across all layers and attention heads, which will be further demonstrated in our Appendix. here likely corresponds to the information necessary for predicting it. To ease visualization and for better comparison, instead of directly showing the attention weights, we rank the tokens and visualize the top 10 tokens with the highest attention weights. For each model, we calculate the token attention ranking for 100 biographies333Because attention paid on meaningless tokens provides little information, we removed periods, commas, spaces, and placeholders at the beginning of a sentence(for example, <bos>)., and summarize the ranking using a bar plot in Figure 2.
@@
Therefore, we can contrast the attention pattern between a large and a small model in the same family to identify the difference in the clue they pay attention to. In Figure 3, we subtract the attention weights of the small model from the large model, and visualize the top 10 tokens with the largest attention differences. The graph shows tokens receiving the most “additional” attention from the large model. It is clear that the correct head entity of the relationship, the “name” tokens (in red color), often receive the most additional attention444The date tokens (in blue color) also appear to rank high in attention differences, which is likely due to the fact that there are on average more date tokens than name tokens in the text, so they are counted more frequently in the top 10 tokens. For example, under the LLaMA 3 tokenizer, the name is split into an average of 3.56 tokens, while the date is split into around 7 tokens..
@@
The token with the -th rank (having the -th largest attention difference) will be dropped out with probability . (A graph of the function is shown in Figure 5). The hyperparameter controls how fast the dropout probability increases with the ranking, and controls the maximum dropout probability. The tokens with higher attention differences will have lower dropout probabilities, encouraging the model to focus more on these tokens. Figure 4 illustrates the process of the proposed augmentation method.
@@ Results
Results in Table 2 show that the proposed token-dropout augmentation based on attention difference significantly outperforms other data augmentation methods. We report QA accuracy on the “university” and the “company” fields as the models have poor performance on these fields under plain finetuning (Table 1). We report exact match (EM) accuracy and normalized word-level F1 scores. We can see that while random dropout and dropout by attention improve performance over no data augmentation, our method achieves much more significant improvement. This proves that contrasting attention of large and small language models indeed finds important but elusive clues in text effectively, and amplifying these clues in the input has immediate positive effects on the model’s memorization efficiency even for the largest 70B model.
@@
Aside from the biography dataset, we also evaluate the proposed method on Wikipedia text to verify if the method helps knowledge learning on general text. Specifically, we evaluate on the Paragraph-Level Wikipedia Question-Answering dataset (Du and Cardie 2018). We first perform continual pretraining on the Wikipedia text paragraphs (included in the dataset), then evaluate the model’s performance on the question-answering data555This is the “closed-book” setting where the model is not allowed to look at the original Wikipedia passage during question answering. It tests the model’s ability to memorize factual knowledge during the continual pretraining phase.. The questions are specifically designed to incorporate coreference dependencies that span multiple sentences in a paragraph, making it a challenging task that tests the model’s ability to learn and memorize complex factual associations.
An example paragraph of Wikipedia text from the dataset is as follows:
@@
| Random token-dropout | 0.7 | - | 0.12 | 0.2233 |
@@
Results in Table 3 show that the proposed method also improves knowledge learning from the Wikipedia text. Naive data augmentation can negatively affect the model’s performance, while our method improves the model’s memorization efficiency by selectively amplifying difficult and elusive clues. This shows that enhancing the model’s focus on important but elusive information in a crucial factor in improving the model’s knowledge learning efficiency in pretraining, and our method is generally applicable to different kinds of text.
@@ Conclusion
Appendix A Data details
Synthetic dataset
The synthetic biography dataset is proposed and used in Zhu and Li (2023). The format of data is similar to Wikipedia passages. It contains multiple forms of information such as birth date, birth place, university and major. The dataset was not directly published by its original authors, and we reconstructed it based on the details and descriptions in their work. The dataset is also used by various related works, such as Khalifa et al. (2024), Golovneva et al. (2024), and Saito et al. (2024).
We used GPT-4 (OpenAI 2023) to generate the biography dataset. The prompt used for generating the dataset is as follows:
Please generate biographies for 100 individuals in the following format:
[name] was born on [month] [date], [year]. He/She spent his/her early years in [city], [country]. He/She received mentorship and guidance from faculty members at [university]. He/She completed his/her education with a focus on [major]. He/She had a professional role at [company].
Below are 2 sentences that have been written to standard, which you can refer to for the format:
“Adrian Wallace was born on March 15, 1975. He spent his early years in Toronto, Canada. He received mentorship and guidance from faculty members at McGill University. He completed his education with a focus on Artificial Intelligence. He had a professional role at the Pacific Tech Innovation Hub.”
“Sofia Ramirez was born on June 22, 1982. She spent her early years in Buenos Aires, Argentina. She received mentorship and guidance from faculty members at the University of São Paulo. She completed her education with a focus on Environmental Science and Policy. She had a professional role at the Green Earth Strategies Corp.”
Please consider the following requirements when generating the biographies:
1. Avoid using existing people’s names, especially those that are well-known.
2. Use real existing places, dates, companies, and majors in the generated biographies.
3. The names should not be repeated.
4. The birth places, universities, and companies should not be strongly associated (e.g., in a same city) for each individual.
For evaluating model’s performance, we use the following question answering task to ask about the 5 pieces of information in the biography about each individual:
Question: When was Liam Thompson born?
Answer: January 5, 1990.
Question: Where was Liam Thompson born?
Answer: Melbourne, Australia.
Question: Which university did Liam Thompson graduate from?
Answer: Sorbonne University.
Question: What did Liam Thompson study?
Answer: Biomedical Engineering.
Question: Where did Liam Thompson work?
Answer: the British Museum.
We use 5-shot prompt in the question answering task, where we provide the model with 5 randomly selected question-and-answer examples of the same format but for well-known individuals. This is to ensure that the model understands the task and the format of the answers. We evaluate the model’s performance using exact match (EM) accuracy and normalized word-level F1 score, used in the SQuAD question answering benchmark (Rajpurkar et al. 2016).
Appendix B Implementation details and discussion
Definition of Distance
“Distance” in Section 3.2 and Figure 1 refers to the number of tokens between the tail entity and the head entity in text, for example, between the “company” field and the person’s name in the biography. The distance is calculated by subtracting the average position of the tokens in the tail entity and the average position of the head entity. In Figure 1, the x-axis marks the information fields with increasing distance. The figure demonstrates a pattern whereby the further away relevant information resides from the name tokens, the lower the model’s accuracy in assimilating this information.
Non-obvious but Important Clues
In the synthetic biography text, the head entity is an important clue and obvious to humans, but is not obvious to language models as seen in their low attention to the head entity (shown in Figure 2). It is not easy for language models to notice the head entity likely due to long distance and the presence of other distractions. We noticed that even large language models can only pay limited attention on those non-obvious but important clues, which provides evidence for the performance improvement of large language models such as LLaMA 3 70B on both synthetic and real-world datasets.
Generalization
The results in Table 2 and Table 3 have demonstrated that our approach can be generalised across different forms of the dataset. And we will show that our method is not confined on certain positions (for example, the end of the sentence) or certain tokens (for example, prepositions).
While we visualize attention at preposition tokens for better illustrating the idea, we show that using attention differences on all token can help enhance knowledge learning (on Wikipedia data) without depending on prepositions. For the biography data, when using attention differences on information fields instead of just on prepositions, the model achieved similar performance improvements with our method on LLaMA 2 7B in Table 4.
| University | Company | |||
| EM | F1 | EM | F1 | |
| LLaMA 2 7B | ||||
| Baselines | ||||
| Plain finetuning | 0.21 | 0.53 | 0.10 | 0.15 |
| Random token-dropout | 0.37 | 0.66 | 0.34 | 0.39 |
| Token-dropout by attention | 0.37 | 0.64 | 0.21 | 0.26 |
| Ours | ||||
| Dropout by all tokens attention diff | 0.53 | 0.75 | 0.51 | 0.59 |
| Dropout by preposition tokens attention diff | 0.67 | 0.84 | 0.56 | 0.59 |
Computational Costs
Since our approach uses a large-size language model and increases the size of data during data augmentation, the cost of models training may increased accordingly. However, we found that the cost increase is often mild in practice.
Firstly, when training a smaller language model, the larger language model is only used for inference and accounts for a small fraction of the computational costs compared to training. Secondly, the increase in training cost is manageable because the model requires significantly fewer epochs to converge with data augmentation. For example, LLaMA 2 70B converges at the 20th epoch without data augmentation, but only at the 6th epoch with data augmentation, as shown in Table 5.
| Epoch | Steps | EM | F1 | |
| LLaMA 2 70B | ||||
| Plain finetuning | ||||
| No Augmentation | 20 | 2000 | 0.58 | 0.62 |
| Ours | ||||
| With Augmentation | 3 | 3300 | 0.60 | 0.72 |
| With Augmentation | 6 | 6600 | 0.86 | 0.89 |
Appendix C More results
Results for other models
| Hyper- | QA performance | |||||
| parameters | University | Company | ||||
| EM | F1 | EM | F1 | |||
| Gemma 2B | ||||||
| Baselines | ||||||
| Plain finetuning | - | - | 0.04 | 0.44 | 0.04 | 0.05 |
| Random token-dropout | 0.6 | - | 0.08 | 0.47 | 0.08 | 0.10 |
| Token-dropout by attention | 0.7 | 0.05 | 0.13 | 0.51 | 0.07 | 0.10 |
| Token-dropout by distance | 0.7 | 0.1 | 0.17 | 0.53 | 0.19 | 0.24 |
| Ours | ||||||
| Dropout by all tokens attention diff | 0.6 | 0.05 | 0.16 | 0.54 | 0.15 | 0.17 |
| Dropout by preposition attention diff | 0.6 | 0.05 | 0.26 | 0.57 | 0.21 | 0.24 |
| LLaMA 2 7B | ||||||
| Baselines | ||||||
| Plain finetuning | - | - | 0.21 | 0.53 | 0.10 | 0.15 |
| Random token-dropout | 0.7 | - | 0.37 | 0.66 | 0.34 | 0.39 |
| Token-dropout by attention | 0.7 | 0.05 | 0.37 | 0.64 | 0.21 | 0.26 |
| Token-dropout by distance | 0.7 | 0.1 | 0.44 | 0.69 | 0.49 | 0.59 |
| Ours | ||||||
| Dropout by all tokens attention diff | 0.6 | 0.05 | 0.53 | 0.75 | 0.51 | 0.59 |
| Dropout by preposition attention diff | 0.7 | 0.05 | 0.67 | 0.84 | 0.56 | 0.59 |
| LLaMA 2 70B | ||||||
| Baselines | ||||||
| Plain finetuning | - | - | 0.48 | 0.71 | 0.58 | 0.62 |
| Random token-dropout | 0.7 | - | 0.71 | 0.86 | 0.68 | 0.72 |
| Token-dropout by attention | 0.7 | 0.05 | 0.70 | 0.83 | 0.77 | 0.80 |
| Token-dropout by distance | 0.7 | 0.1 | 0.79 | 0.89 | 0.87 | 0.95 |
| Ours | ||||||
| Dropout by all tokens attention diff | 0.6 | 0.05 | 0.75 | 0.88 | 0.71 | 0.80 |
| Dropout by preposition attention diff | 0.7 | 0.05 | 0.85 | 0.92 | 0.89 | 0.92 |
We have done experiments on LLaMA 2 and Gemma models on both the Synthetic dataset and Real-world dataset. Results are shown in Table 6 and Table 8, which strengthen the persuasiveness of our approach and demonstrate that the method can be generalised across different models and data. We introduced another two methods (distance dropout and all attention diff) in our experiments for illustrating the effectiveness and generality of our approach.
In the original synthetic biography dataset, the head entity (the person’s name) appears at the beginning of the sentence. For more general scenarios, we also consider a version of the biography dataset, where the head entity (the person’s name) appears in a random sentence within the paragraph. For example:
He was born on January 5, 1990. He spent his early years in Melbourne, Australia. Liam Thompson received mentorship and guidance from faculty members at Sorbonne University. He completed his education with a focus on Biomedical Engineering. He had a professional role at the British Museum.
We also implement a distance-based augmentation baseline that dropout tokens based on the distance between the head entity and the tail entity.
We test our method on this dataset and show that our method also outperforms the distance-based augmentation method, showing that our method can adaptively enhance clues on a random position based on its semantic significance rather than simply enhancing clues that are far-away. The results are listed in Table 7 (LLaMA 2):
| University | Company | |||
| EM | F1 | EM | F1 | |
| LLaMA 2 7B | ||||
| Baselines | ||||
| Plain finetuning | 0.14 | 0.47 | 0.27 | 0.29 |
| Random token-dropout | 0.21 | 0.55 | 0.29 | 0.31 |
| Token-dropout by attention | 0.16 | 0.24 | 0.30 | 0.58 |
| Token-dropout by distance | 0.18 | 0.54 | 0.41 | 0.45 |
| Ours | ||||
| Dropout by all tokens attention diff | 0.34 | 0.64 | 0.62 | 0.66 |
| Hyperparameters | QA performance | |||
| EM | F1 | |||
| Gemma 2B | ||||
| Baselines | ||||
| Plain finetuning | - | - | 0.105 | 0.186 |
| Random token-dropout | 0.7 | - | 0.101 | 0.189 |
| Token-dropout by attention | 0.7 | 0.05 | 0.098 | 0.193 |
| Ours | ||||
| Dropout by preposition attention diff | 0.7 | 0.005 | 0.115 | 0.204 |
| LLaMA 2 7B | ||||
| Baselines | ||||
| Plain finetuning | - | - | 0.138 | 0.231 |
| Random token-dropout | 0.7 | - | 0.126 | 0.217 |
| Token-dropout by attention | 0.7 | 0.05 | 0.123 | 0.217 |
| Ours | ||||
| Dropout by preposition attention diff | 0.7 | 0.03 | 0.144 | 0.268 |
| LLaMA 2 70B | ||||
| Baselines | ||||
| Plain finetuning | - | - | 0.179 | 0.282 |
| Random token-dropout | 0.7 | - | 0.187 | 0.307 |
| Token-dropout by attention | 0.7 | 0.05 | 0.204 | 0.321 |
| Ours | ||||
| Dropout by preposition attention diff | 0.7 | 0.005 | 0.219 | 0.331 |
Attention visualization
Based on the principle of simplicity, we show that a simple averaging of attention weights across all layers and heads already demonstrated the main idea of the paper and produces substantial improvements. We believe that analysis of individual layers and heads may lead to more specific and better results, and we leave this for future work.

### v2

@@ Abstract
Code — https://github.com/tsinghua-msiip/contrasting˙attention
@@ Introduction
We release the code and data used in this paper for reproducibility and further research.
@@
The questions are formatted like “When was Liam Thompson born?”. When questions were rephrased using GPT-4, performance generally declined, indicating that the original questioning format yielded the best performance, so that question style has minimal impact on our conclusion.
@@
We show that even pretrained models with billions of parameters struggle to memorize facts perfectly in continual pretraining. Table 1 show that while Gemma 2 (Team et al. 2024) and LLaMA 3 (Dubey et al. 2024) memorize the first two pieces of information (birth date and birth city) with high accuracy, they struggle to memorize the following three pieces of information (university, major, and company). This rules out the possibility that the performance deficiency is due to limited model size or insufficient pretraining. We also tried swapping the positions of five kinds of information resulted in the same trend: accuracy decreases as distance increases, demonstrating that long-range dependencies, rather than en-tity types, are the primary cause of poor performance.
@@
| Date | City | University | Major | Company | ||
@@
We have shown that language models could achieve near-perfect accuracy in memorizing relationships that span a short distance in text, but struggle when they span a longer distance. In this section, we use attention weights as an interpretability tool to analyze the model’s behavior while learning long-range dependencies. We show that LLMs can pay inadequately little attention to key information that is located further away, and more performant larger models can pay more attention to these information than smaller models.
@@
To visualize model’s attention weights when predicting the tail entities in a relationship, we extract the attention weights at the preposition tokens, i.e., the word immediately preceding the tail entity. For example, in the sentence “He received mentorship and guidance from faculty members at Sorbonne University”, the attention from the token “at” is extracted. Because the model is predicting the tail entity “Sorbonne University” at this position, the attention weights111To simplify analysis, we took the approach of averaging the attention weights across all layers and attention heads. here likely corresponds to the information necessary for predicting it. To ease visualization and for better comparison, instead of directly showing the attention weights, we rank the tokens and visualize the top 10 tokens with the highest attention weights. For each model, we calculate the token attention ranking for 100 biographies222Because attention paid on meaningless tokens provides little information, we removed periods, commas, spaces, and placeholders at the beginning of a sentence(for example, <bos>)., and summarize the ranking using a bar plot in Figure 2.
@@
Therefore, we can contrast the attention pattern between a large and a small model in the same family to identify the difference in the clue they pay attention to. In Figure 3, we subtract the attention weights of the small model from the large model, and visualize the top 10 tokens with the largest attention differences. The graph shows tokens receiving the most “additional” attention from the large model. It is clear that the correct head entity of the relationship, the “name” tokens (in red color), often receive the most additional attention333The date tokens also appear to rank high in attention differences, which is likely due to the fact that there are on average more date tokens than name tokens in the text, so they are counted more frequently in the top 10 tokens. For example, under the LLaMA 3 tokenizer, the name is split into an average of 3.56 tokens, while the date is split into around 7 tokens..
@@
The token with the -th rank (having the -th largest attention difference) will be dropped out with probability . The hyperparameter controls how fast the dropout probability increases with the ranking, and controls the maximum dropout probability. The tokens with higher attention differences will have lower dropout probabilities, encouraging the model to focus more on these tokens. Figure 4 illustrates the process of the proposed augmentation method.
@@ Results
Results in Table 2 show that the proposed token-dropout augmentation based on attention difference significantly outperforms other data augmentation methods. We report QA accuracy on the “university” and the “company” fields as the models have poor performance on these fields under plain finetuning (Table 1). We report exact match (EM) accuracy and normalized word-level F1 scores. We can see that while random dropout and dropout by attention improve performance over no data augmentation, our method achieves much more significant improvement. We also collected the results of other information from models trained in our method and accuracy increased across models. This proves that contrasting attention of large and small language models indeed finds important but elusive clues in text effectively, and amplifying these clues in the input has immediate positive effects on the model’s memorization efficiency even for the 70B model.
@@
Aside from the biography dataset, we also evaluate the proposed method on Wikipedia text to verify if the method helps knowledge learning on general text. Specifically, we evaluate on the Paragraph-Level Wikipedia Question-Answering dataset (Du and Cardie 2018). We first perform continual pretraining on the Wikipedia text paragraphs (included in the dataset), then evaluate the model’s performance on the question-answering data444This is the “closed-book” setting where the model is not allowed to look at the original Wikipedia passage during question answering. It tests the model’s ability to memorize factual knowledge during the continual pretraining phase.. The questions are specifically designed to incorporate coreference dependencies that span multiple sentences in a paragraph, making it a challenging task that tests the model’s ability to learn and memorize complex factual associations.
An example of Wikipedia text from the dataset is:
@@
| Random token-dropout | 0.7 | - | 0.12 | 0.223 |
@@
Results in Table 3 show that the proposed method also improves knowledge learning from the Wikipedia text. Unlike naive data augmentation, our method improves the model’s memorization efficiency by selectively amplifying difficult and elusive clues. This shows that enhancing the model’s focus on important but elusive information is a crucial factor in improving knowledge learning efficiency, and our method is generally applicable to different kinds of text.
@@ Conclusion
Acknowledgements
This work was supported by the Noncommunicable Chronic Diseases-National Science and Technology Major Project (Grant No. 2023ZD0506501).